# latexYou have requested package

### v1

@@ Abstract
Activation functions introduce non-linearity into Neural Networks, enabling them to learn complex patterns. Different activation functions vary in speed and accuracy, ranging from faster but less accurate options like ReLU to slower but more accurate functions like SiLU or SELU. Typically, same activation function is used throughout an entire model architecture. In this paper, we conduct a comprehensive study on the effects of using mixed activation functions in YOLO-based models, evaluating their impact on latency, memory usage, and accuracy across CPU, NPU, and GPU edge devices. We also propose a novel approach that leverages Neural Architecture Search (NAS) to design YOLO models with optimized mixed activation functions.The best model generated through this method demonstrates a slight improvement in mean Average Precision (mAP) compared to baseline model (SiLU), while it is 22.28% faster and consumes 64.15% less memory on the reference NPU device.
@@ 1 Introduction
Improving the accuracy of computer vision models has become a highly competitive area, with researchers constantly refining or developing new architectures to achieve state-of-the-art performance. For instance, the recent YOLO10 [5] architecture has set a new benchmark for object detection within the YOLO family [23], outperforming previous models on the popular Microsoft COCO dataset [14]. The current trend toward higher accuracy often involves designing more complex models by increasing the number of learnable parameters or adding performance-enhancing blocks. However, these gains in accuracy typically come at the cost of higher latency. A common practice in deep learning model design is to use a single activation function across all layers. Activation functions introduce non-linearity, enabling models to learn complex patterns from data. Since these functions are applied throughout the architecture, they significantly influence both model latency and accuracy. For example, the Sigmoid Linear Unit (SiLU) [22] offers better accuracy but operates slower than the faster, but less accurate, Rectified Linear Unit (ReLU)[2]. Through our exploration, we observed that employing different activation functions for various layers in the model can improve mean Average Precision (mAP) by 1-2% and reduce latency by 20-30%, compared to the baseline model using SiLU activation across all layers. In this study, we controlled all other factors that could affect latency and accuracy to isolate the impact of mixed activation function design. Additionally, we propose a Neural Architecture Search (NAS) approach to automatically select the optimal activation function for each layer based on given latency and accuracy constraints.
We summarize our contributions as follows:
  • We introduce ActNAS, a NAS based method to search for the best configuration of the model using mixed activation function for each layer.
@@
  • We demonstrate that our ActNAS generated models have better mAP and lower latency compared to baseline models trained on COCO dataset and tested on difference edge devices.
@@ 2 Related Works
In building computer vision models, it is common practice to use the same activation function across all layers of the architecture. For example, YOLO models[23], whether in the n, m, s, l, or tiny variants, consistently use the SiLU [22] activation function due to its superior accuracy compared to alternatives like ReLU [19]. However, this higher accuracy often comes at the cost of increased latency, which opens up the possibility of exploring other activation functions such as ReLU[2], LeakyReLU[7], or Hardswish[3]. While YOLO models do allow the option to switch from SiLU to other activations across the entire architecture, there has been no prior research—based on our knowledge—on using mixed activation functions within the same YOLO model. The closest existing work involves designing hybrid activation functions, simplifying blocks in the architecture by applying a different activation function than the main one or mixing activations in image classification models.
@@ 3 Methodology
Unlike the straightforward approach of using the same activation function across all layers in a model, mixed activation models incorporate multiple activation functions within the architecture. As illustrated in Figure 4, the YOLO5n model with SiLU activation achieves the highest mAP but is the slowest across different devices, while the model with ReLU activation is the fastest, though it experiences a significant drop in mAP. The goal of mixed activation models is to identify the optimal combination of activation functions for each layer, creating a model that strikes the best balance between latency and accuracy.
For our experiments, we use the YOLO5n model combined with the YOLO8n head (which offers better accuracy), with the SiLU-based model serving as the reference. As shown in Figure 2, we generate a search space by systematically replacing each activation function in the reference model with a set of candidate activations: [ReLU, SiLU, Hardswish, ReLU6, and LeakyReLU]. We replace one activation at a time, as depicted in Figure 1, resulting in 345 candidate models for YOLO5n (since the reference model has 69 activations and there are 5 candidate activations for each layer).
To evaluate the impact of each activation replacement, we create a new model for each candidate and measure its accuracy, latency, and memory usage, comparing it to the reference model. This process is iterative and is repeated for all 345 candidate models. The results are recorded in performance tables: the latency table, accuracy table, and memory table. Each entry in the accuracy table includes the layer name, activation name, reference accuracy, and delta accuracy (the difference between the reference model’s accuracy and the accuracy of the model with the replaced activation). Similarly, the latency and memory tables follow the same structure but record latency and memory values instead of accuracy.
@@ 3.1 Accuracy Estimator
As the number of activation and layer combinations increases, the search space for mixed activation models grows exponentially, making it computationally expensive and time-consuming to train each model and evaluate the impact of replacing activations. To address this, we use the NWOT Zero-Cost (ZC) metric [16] to estimate the effect of changing activation functions in individual layers on model accuracy. We validated that the NWOT score correlates strongly with the accuracy of fully trained models using SiLU, ReLU, and Hardswish activations. This high correlation indicates that the NWOT score can reliably predict the impact of per-layer activation changes on accuracy without the need for full training. For each candidate activation replacement, the NWOT score is calculated, and the accuracy table is updated accordingly, as described in the previous section. This approach significantly reduces the computational cost and time required to explore the mixed activation model search space.
Figure 2: Model Benchmarking using training free estimators and on device inference
@@ 3.2 Latency Computation
The impact of different activation functions on latency and memory depends on the runtime and hardware architecture. For this reason, a variety of hardware, including two CPUs (ARM Cortex-A57 in Jetson Nano and ARM Cortex-A53 on Raspberry Pi 3), an embedded GPU (Jetson Nano), and a reference Neural Processing Unit (NPU) is used for our experiments. To measure the performance of the models on these devices, ONNX Runtime [18] for Jetson Nano, TensorFlow Lite (TFLite) [15] for the ARM Cortex processors, and a custom compiler/runtime for the NPU is used. Each model is converted to ONNX, TFLite, or a custom format depending on the hardware. Latency values are recorded as average over 50 runs to ensure correctness. The input size of 224x224 is used for NPU due to memory limitations while other latency numbers are calculated using 640x640 input. This variation in hardware and runtime ensures that the results aren’t biased towards a specific platform and or runtime, providing a balanced perspective of how different activations perform across different systems.
@@ 3.3 Mixed Activation Model Search
Latency, memory and accuracy table obtained from benchmarking step is used to construct mixed activation models.
@@ 3.3.1 Local Zero Cost Maxima Approach
Our simple approach to create a model using just two activation functions, such as SiLU and ReLU. Starting with a reference model that uses only SiLU, we apply the Local Zero Cost Maxima (LZCM) method to iteratively replace activations with ReLU based on the zero-cost NWOT score. If the NWOT score improves when a specific layer’s SiLU activation is replaced with ReLU, that replacement is kept. The modified model is then trained from scratch on the COCO dataset. Using this simple approach, we developed two models: LZCM1, which mixes SiLU and ReLU, and LZCM2, which mixes Hardswish and ReLU activations.
@@ 3.3.2 Naive Approach
Benchmarking models generated using the LZCM approach shows that activations in layers closer to the input have a significant impact on both latency and memory due to the larger feature sizes in those layers. Replacing the activations in the initial layers with ReLU significantly reduces latency and memory usage. In our simple approach, we replace the first three activations near the input with ReLU, while keeping the remaining layers as SiLU. This leads to an efficient balance between performance and resource consumption.
@@ 3.3.3 Activation NAS Approach
The LZCM method follows a straightforward decision-making process, replacing activations in individual layers based on their zero-cost score. However, this approach only optimizes each layer locally and doesn’t account for the combined impact of replacing activations in multiple layers at once. In contrast, the Activation NAS (ActNAS) approach searches for the optimal activation function for each layer while considering a global budget for latency, memory, or accuracy. The search space for ActNAS is defined by the number of layers multiplied by the number of activation function options. To navigate this space, we use the latency, memory, and accuracy tables, along with predefined constraints, to guide the search. For ActNAS, we experimented with both random search and Integer Linear Programming (ILP) to find the best activation configurations.
@@
As shown in Figure 3, generating model proposals using Activation NAS requires pre-computed latency and accuracy tables. These tables capture the changes in these metrics when a single layer’s activation in the original YOLO model (with SiLU activation) is replaced by an alternative activation. The tables are then converted into latency and accuracy cost matrices, where the rows represent the layers being modified and the columns represent candidate activations (SiLU, ReLU, Hardswish, LeakyReLU, and ReLU6). In addition to these matrices, we also define latency or accuracy constraints, which serve as the overall budget that the algorithm can work within when constructing new models with mixed activations. The random search process begins by randomly selecting activations from the search space that fit within the given constraints. It then computes the total cost (in terms of latency or accuracy) for the model and marks it as the best possible model at that point. The algorithm repeats the random search, comparing each newly constructed model with the previous best. If a new model has a better overall cost, it updates the best model definition and continues iterating through the random selection process.
ILP Search
To generate model proposals with mixed activations, we first constructed latency and accuracy matrices, where each element represents the difference in metrics from the original SiLU-based YOLO model when a single layer’s activation is replaced. We then formulated this problem as an Integer Linear Programming (ILP) optimization problem. The objective was to minimize the latency and accuracy cost while adhering to constraints on the overall latency or accuracy of the newly constructed models. In the ILP formulation, one variable vector represents the indices of the layers where activations are applied, and the other represents the candidate activations for those layers. We used the open-source PuLP library [17], a Python toolkit for linear programming, to solve this ILP problem. The result is a ranked list of the top k model proposals. To ensure diversity among proposals, the solution avoids excessive overlap, keeping the model architectures meaningfully different.
In Tables 2 and 1, the ActNAS1 model was constructed by applying an accuracy constraint while minimizing latency for the Cortex A-53 CPU hardware. The ILP search can be adapted for different hardware targets by imposing a latency constraint on the constructed model. For example, ActNAS2 and ActNAS3 were built by setting constraints for latency on the Jetson Nano GPU and Cortex A-53 CPU, respectively, with the goal of minimizing accuracy cost while staying within the latency budget for each device.
3.4 Model Fine-tuning
We used Ultralytics [11] to train and fine-tune all our models. For consistency across experiments, we kept all hyperparameters exactly the same, training each model from scratch for 300 epochs using $640\times 640640 × 640$ COCO images and evaluating them on the COCO validation set. While mixed activation models could be fine-tuned for a few epochs starting from pre-trained weights, we chose to train all models from scratch to ensure a fair comparison.
Figure 3: Activation NAS Process
Table 1: Performance of ActNAS models compared to baseline models on GPU and CPU
@@ 3.4 Model Fine-tuning
YOLO5n_SiLU 0.3400 27.20 - - 937.89 - -
YOLO5n_ReLU 0.3205 20.18 - - 793.27 - -
YOLO5n_Hardswish 0.3342 25.63 - - 823.91 - -
LZCM1(SiLU/ReLU) 0.3360 26.61 -21.54% 2.17% 916.97 -0.58% 2.23%
LZCM2(SilU/Hswish) 0.3346 31.15 -21.54% -14.52% 828.72 -0.58% 11.64%
Naive(ReLU/SiLU) 0.3380 24.55 4.21% 9.74% 906.78 -10.06% 3.32%
ActNAS1(Mixed) 0.3420 24.78 3.32% 8.90% 809.96 1.69% 13.64%
ActNAS2(Mixed) 0.3320 23.50 8.31% 13.60% 966.74 -17.34% -3.08%
ActNAS3(Mixed) 0.3400 24.51 4.37% 9.89% 965.64 -17.20% -2.96%
@@
Model mAP NPU NPU
Latency Improvement (%) RAM (KB) Improvement (%)
(ms) Hswish SiLU Hswish SiLU
YOLO5n_SiLU 0.3400 22.35 - - 1230.00 - -
YOLO5n_ReLU 0.3205 17.46 - - 588.00 - -
YOLO5n_Hardswish 0.3342 18.53 - - 392.00 - -
LZCM1(SiLU/ReLU) 0.3360 21.87 -5.50% 2.15% 1200.00 -206.12% 2.44%
LZCM2(SiLU/Hswish) 0.3346 19.55 -5.50% 12.53% 624.75 -59.38% 49.21%
Naive(SiLU/ReLU) 0.3380 21.43 -15.65% 4.12% 661.50 -68.75% 46.22%
ActNAS1(Mixed) 0.3420 17.37 6.26% 22.28% 514.50 -31.25% 58.17%
ActNAS2(Mixed) 0.3320 19.27 -3.99% 13.78% 520.63 -32.81% 57.67%
ActNAS3(Mixed) 0.3400 17.50 5.56% 21.70% 441.00 -12.50% 64.15%

### v2

# ActNAS : Generating Efficient YOLO Models using Activation NAS

@@ Abstract
Activation functions introduce non-linearity into neural networks, allowing them to learn complex patterns. Different activation functions impact both speed and accuracy; for instance, ReLU is fast but often less precise, while SiLU offers higher accuracy at the expense of speed. Traditionally, a single activation function is used across the entire model. In this work, we conducted a comprehensive study on the effects of using mixed activation functions in YOLO-based models, examining their impact on latency, memory usage, and accuracy across CPU, NPU, and GPU edge devices. We propose a novel approach, Activation NAS (ActNAS), which utilizes Hardware-Aware Neural Architecture Search (HA-NAS) to design YOLO models with mixed activation functions optimized for specific hardware. Models generated through ActNAS achieve similar mean Average Precision (mAP) compared to baseline models, while running up to 1.67 times faster and/or using 64.15% less memory on target devices. Additionally, we demonstrate that hardware-aware models learn to leverage architectural and compiler-level optimizations, resulting in highly efficient performance tailored to each hardware platform.
@@ 1 Introduction
Improving the accuracy of computer vision models has become a highly competitive area, with researchers constantly refining or developing new architectures to achieve state-of-the-art performance. For example, the recent YOLO10 [25] architecture has set a new benchmark for object detection within the YOLO family [23], outperforming previous models in the popular Microsoft COCO dataset [12]. The current trend toward higher accuracy often involves designing more complex models by increasing the number of learnable parameters or adding performance-enhancing blocks. However, these gains in accuracy typically come at the cost of higher latency. Some research works redesign activation functions to learn complex patterns and improve accuracy-latency tradeoffs. Since these activation functions are applied throughout the model architecture, they significantly influence both model latency and accuracy. For example, the Sigmoid Linear Unit (SiLU) [4] offers better accuracy at the expense of higher latency (slower) than the less accurate but less latent (faster) Rectified Linear Unit (ReLU) [2] and HardSwish [7] activation function offering the best latency-accuracy tradeoff. Figure 1 demonstrates the accuracy-latency tradeoff of these variations for YOLO5n and YOLO8m models. A common practice in deep learning model design is using a single activation function across all layers in the model architecture. Through our exploration, we observed that employing different activation functions for various layers in the model either improves the mean Average Precision (mAP) by less than 1% or stays within 1% drop but reduces latency by 30-70%, compared to the baseline model using SiLU activation across all layers. In this study, we controlled all other factors that could affect latency and mean average precision (mAP) to isolate the impact of the mixed activation function design. We also propose a Hardware Aware Neural Architecture Search (HA-NAS) approach to automatically select the optimal activation function for each layer based on given latency and accuracy constraints.
Figure 1: Latency-mAP plot of YOLO5n and YOLO8m for two different NPUs. SiLU models are slowest and ReLU models are fastest on both devices
We summarize our contributions as follows.
  • We introduce ActNAS, a Hardware Aware NAS based method to search for the best configuration of the model using a mixed activation function for each layer.
@@
  • We demonstrate that our ActNAS generated models have negligible drop in mAP, but significant improvement over latency compared to the baseline models trained on COCO dataset and tested on difference edge devices.
  • We also demonstrate the hardware aware model for one device may not necessarily perform well on other type of device.
@@ 2 Related Works
In building computer vision models, it is a common practice to use the same activation function across all layers of the architecture. For example, YOLO models [23], whether in the n, m, s, l, or tiny variants, consistently use the SiLU [4] activation function due to its superior accuracy compared to alternatives like ReLU [21]. However, this higher accuracy often comes at the cost of increased latency, which opens up the possibility of exploring other activation functions such as ReLU [2], LeakyReLU[26], or Hardswish [7]. While YOLO models do allow the option to switch from SiLU to other activations across the entire architecture, there has been no prior research (based on our knowledge) on using mixed activation functions within the same YOLO model. The closest existing work involves designing hybrid activation functions, simplifying blocks in the architecture by applying a different activation function than the main one or mixing activations in image classification models.
@@ 3 Methodology
Unlike the straightforward approach of using the same activation function across all layers in a model, mixed activation models incorporate multiple activation functions within the architecture. As illustrated in Figure 1, the YOLO5n and YOLO8m models with SiLU activation achieve higher mAP but is least efficient on different devices, while the model with ReLU activation is the fastest with lower mAP. The goal of mixed activation models is to identify the optimal combination of activation functions for each layer, creating a model that strikes the best balance between latency and accuracy.
Figure 3: mAP vs NWOT cost for 131 MCUbench[22] Models. Each point represents one model and NWOT and mAP score of each model is plotted in orange and blue colors respectively
For our experiments, we used YOLO5n combined with YOLO8 head (which offers better accuracy) and YOLO8m model, with the SiLU-based model serving as a reference. As shown in Figure 4, we generate a search space by systematically replacing each activation function in the reference model with a set of candidate activations: [ReLU, SiLU, Hardswish, ReLU6, and LeakyReLU]. We replace one activation at a time, as depicted in Figure 2, resulting in 345 candidate models for YOLO5n (since the reference model has 69 activations and there are 5 candidate activations for each layer).
To evaluate the impact of each activation replacement, we create a new model for each candidate and measure its accuracy, latency, and memory usage, comparing it to the reference model. This process is iterative and is repeated for all 345 candidate models. The results are recorded in performance tables: the latency table, accuracy table, and memory table. Each entry in the accuracy table includes the layer name, activation name, reference accuracy, and delta accuracy (the difference between the reference model’s accuracy and the candidate model). The latency and memory tables follow the same process and record latency and memory values instead of accuracy.
@@ 3.1 Accuracy Estimator
As the number of activations and layer combinations increase, the search space grows exponentially, making it computationally expensive and time-consuming to train each model and evaluate the impact of replacing activations. To address this, we use the NWOT Zero-Cost (ZC) metric [15] to approximate the impact of changing activation functions in different layers on the accuracy of the model.
To validate the effectiveness of NWOT score as a proxy to the mAP of trained YOLO models, we estimated the correlation between the mAP of 131 variations of YOLO models in MCUBench [22] and their corresponding NWOT scores. Figure 3 shows that the NWOT score has a strong correlation with the mAP of fully trained models. The high correlation indicates that the NWOT score can reliably predict the impact of per-layer activation changes on accuracy without the need of training each candidate model. For each candidate activation replacement, the NWOT score is calculated, and the accuracy table is updated accordingly as described in the previous section. This approach significantly reduces the computational cost and time required to explore the large search space of mixed activation models. The suitability of zero cost proxy based method also opens an avenues for more complex NAS methods requiring larger search space.
Figure 4: Model Benchmarking using training free estimators and on device inference. Each candidate model is created by replacing just one activation function.
@@ 3.2 Latency Computation
The impact of different activation functions on latency and memory depends on the hardware architecture, compiler level optimization and the runtime. For this reason, a variety of hardware, including two CPUs (ARM Cortex-A57 on Jetson Nano and ARM Cortex-A53 on Raspberry Pi 3), an embedded GPU (Jetson Nano), and three different reference neural processing units (NPU1, NPU2, NPU3) are used for our experiments. We are not disclosing the vendor name and actual name of these NPUs as some of these are yet to available in the market. To measure the performance of the models on these devices, we used ONNX Runtime [18] for Jetson Nano, TensorFlow Lite (TFLite) [13] for ARM Cortex processors, and custom compilers/runtimes as needed by each NPU. Each model is converted to ONNX, TFLite, or a custom format depending on the compiler. Latency values are recorded as average over 50 runs to ensure correctness. The input size of $224\times 224224 × 224$ or $640\times 640640 × 640$ is used depending upon the memory limitations. This variation in hardware, runtime and model family ensures that the results aren’t biased towards a specific platform, runtime or model family, providing a balanced perspective of how different activations perform across different systems.
@@ 3.3 Mixed Activation Model Search
Latency, memory and accuracy table obtained from the benchmarking step is used to construct mixed activation models.
@@ 3.3.1 Local Zero Cost Maxima Approach
We started with the simple approach to create a model using just two activation functions, such as SiLU and ReLU. Starting with a reference model that uses only SiLU, we apply the Local Zero Cost Maxima (LZCM) method to iteratively replace activation of a layer with ReLU provided the NWOT score is higher compared to the model that uses SiLU for the same layer. The modified model is then trained from scratch on the COCO dataset. Using this simple approach, we developed two models: LZCM1, by using SiLU and ReLU, and LZCM2, by using Hardswish and ReLU activations.
Figure 5: Activation NAS Process
@@ 3.3.2 Naive Approach
Benchmarking models generated using the LZCM approach show that activations in layers closer to input have a significant impact on both latency and memory due to the larger feature sizes in those layers. Replacement of activations in initial layers with ReLU significantly reduces latency and memory usage. In our simple approach, we replace the first three activations near the input with ReLU, while keeping the remaining layers as SiLU. This leads to an efficient balance between performance and resource consumption.
@@ 3.3.3 Activation NAS Approach
The LZCM method follows a straightforward decision-making process, replacing activations in individual layers according to their zero-cost score. However, this approach only optimizes each layer locally and does not account for the combined impact of replacing activations in multiple layers at once. In contrast, the Activation NAS (ActNAS) approach searches for the optimal activation function for each layer while considering a global budget for latency, memory, or accuracy. The search space for ActNAS is defined by the number of layers multiplied by the number of activation function options. To navigate this space, we used the latency, memory, and accuracy tables, along with predefined constraints, to guide the search. We experimented with both random search and Integer Linear Programming (ILP) [24] to find the best activation configurations.
@@
As shown in Figure 5, generating model proposals using Activation NAS requires precomputed latency and accuracy tables. These tables capture the changes in these metrics when a single layer’s activation in the original YOLO model (with SiLU activation) is replaced by an alternative activation. The tables are then converted into latency and accuracy cost matrices, where the rows represent the layers being modified, and the columns represent candidate activations (SiLU, ReLU, Hardswish, LeakyReLU and ReLU6). In addition to these matrices, we also define latency or accuracy constraints, which serve as the overall budget that the algorithm can work within when constructing new models with mixed activations. The random search process begins by randomly selecting activations from the search space that fit within the given constraints. Then it computes the total cost (in terms of latency or accuracy) for the model and marks it as the best possible model at that point. The algorithm repeats the random search, comparing each newly constructed model with the previous best. If a new model has a better overall cost, it updates the best model definition and continues iterating through the random selection process.
Table 1: Performance of ActNAS models compared to baseline YOLO5n models on GPU and CPU
@@
SiLU 33.60 27.20 - - 937.89 - -
ReLU 31.88 20.18 - - 793.27 - -
Hardswish 33.13 25.63 - - 823.91 - -
LZCM1(SiLU/ReLU) 33.13 26.61 -21.54% 2.17% 916.97 -0.58% 2.23%
LZCM2(SilU/Hswish) 33.29 31.15 -21.54% -14.52% 828.72 -0.58% 11.64%
Naive(ReLU/SiLU) 33.52 24.55 4.21% 9.74% 906.78 -10.06% 3.32%
ActNAS5n.1 32.80 24.78 3.32% 8.90% 809.96 1.69% 13.64%
ActNAS5n.2 32.20 23.50 8.31% 13.60% 966.74 -17.34% -3.08%
@@
Model mAP NPU1 NPU1
Latency Improvement (%) RAM Improvement (%)
(ms) Hswish SiLU (KB) Hswish SiLU
SiLU 0.3366 22.35 - - 1230.00 - -
ReLU 0.3188 17.46 - - 588.00 - -
Hardswish 0.3313 18.53 - - 392.00 - -
LZCM1(SiLU/ReLU) 0.3313 21.87 -5.50% 2.15% 1200.00 -206.12% 2.44%
LZCM2(SiLU/Hswish) 0.3329 19.55 -5.50% 12.53% 624.75 -59.38% 49.21%
Naive(SiLU/ReLU) 0.3352 21.43 -15.65% 4.12% 661.50 -68.75% 46.22%
ActNAS5n.1 0.3280 17.37 6.26% 22.28% 514.50 -31.25% 58.17%
ActNAS5n.2 0.3250 17.50 5.56% 21.70% 441.00 -12.50% 64.15%
ILP Search
To generate model proposals with mixed activations, we first created latency and accuracy matrices, where each element represents the difference in metrics from the original SiLU-based YOLO model when a single-layer activation is replaced. We then formulated this problem as an Integer Linear Programming (ILP) [24] optimization problem. The objective was to minimize the cost of latency and accuracy while adhering to the constraints on the overall latency or accuracy of the newly constructed models. In the ILP formulation, one variable vector represents the indices of the layers where activations are applied, and the other represents the candidate activations for those layers. We used the open-source PuLP library [16], a Python toolkit for linear programming, to solve this ILP problem. The result is a ranking list of the top k model proposals. To ensure diversity among proposals, the solution avoids excessive overlap, keeping the model architectures meaningfully different.
The ActNAS models (ActNAS5n.1 is NPU1 aware and ActNAS5n.2 is Jetson Nano aware) in Table 1 and 2 are hardware aware mixed activation models (YOLO5n) which means that we used the latency of each candidate models on NPU1 to create the optimized models. Similarly, ActNAS8m.1 and ActNAS8m.2 models in Table 3 are NPU2 aware mixed activation models (YOLO8m). The ILP search can be adapted for different hardware targets by imposing a latency constraint on the constructed model.
3.4 Model Fine-tuning
We used Ultralytics [8] to train and fine-tune all our models. For consistency across experiments, we kept all hyperparameters exactly the same, training each model from scratch for 300 epochs using $640\times 640640 × 640$ COCO images and evaluating them on the COCO validation set which contains $50005000$ images. While mixed activation models could be fine-tuned for a few epochs starting from pre-trained weights, we chose to train all models from scratch to ensure a fair comparison.
Table 3: Hardware aware performance of YOLO8m ActNAS models on different NPUs. All latency numbers are for $640\times 640640 × 640$ images. NPU1 is a low end NPU.
Model mAP Latency(ms)
NPU1 NPU2 NPU3
x100
SiLU 48.57 241.50 190.87 340.00
HardSwish 48.96 166.80 140.33 308.80
ReLU 48.03 93.17 103.73 283.40
ActNAS8m.1 48.50 188.40 136.85 319.60
ActNAS8m.2 48.60 143.80 114.33 305.48